OpenStudy (anonymous):

find the limit as x approaches 9 : WILL GIVE EQUATION IN RESPONSE

OpenStudy (anonymous):

\[\lim_{x \rightarrow 9}(\frac{ 1 }{ \sqrt{x}-3 } - \frac{ 6 }{ x-9 })\]

OpenStudy (aravindg):

do you know l hospital's rule?

OpenStudy (anonymous):

please don't use lhopitals or anything im in basic calculus and haven't learned it yet. Suposed to be able to algebriaically change it to solve. I know the answer is suposed to be 1/6 i just don't know how to get there

OpenStudy (aravindg):

ok by basics , rationalise first fraction

OpenStudy (aravindg):

multiply by \(\sqrt{x}+3\) on numerator and denominator

OpenStudy (anonymous):

so the left turns into right?

OpenStudy (anonymous):

ugh didn't get the equation in

OpenStudy (anonymous):

\[\frac{ \sqrt{x}+3 }{ x-3 }\]

OpenStudy (aravindg):

try again ?

OpenStudy (anonymous):

is that not right after multiplying by \[\sqrt{x}\]+3?

OpenStudy (anonymous):

so i now have \[\frac{ \sqrt{x}+3 }{ x-9 } - \frac{ 6 }{ x-9}\]

OpenStudy (aravindg):

ok now add them !! note denominator is same

OpenStudy (anonymous):

well that gives me zero... and i know the answer is suposed to be 1/6

OpenStudy (aravindg):

try again?

OpenStudy (anonymous):

right, nevermind. thanks!

OpenStudy (anonymous):

no i still get zero

OpenStudy (anonymous):

\[\frac{ \sqrt{x}+3 }{ x-9 }- \frac{ 6 }{ x-9 }\]

OpenStudy (aravindg):